OpenStudy (anonymous):

Rewrite the equation in slope-intercept form. Then identify the slope and the y-intercept of the line. 1.x-y=6 2.-12x-4y=2 3.-x-10y=20

OpenStudy (anonymous):

1. y =x-6 , 2. y=3x+(1/6) , 3. y=x/10 +2

OpenStudy (anonymous):

1. y= x-6 2. -3x-1 3. y= -x/10- 2

OpenStudy (anonymous):

number 2 from raheen is incorrect

OpenStudy (anonymous):

so whats the yintercept and slope for each one.